OpenStudy (j.f):

the graph of a straight line y=f(x) passes through the point (3,5) and (4,a). If the slope of the line is 3, what is the value of a? I'm drawing a blank as to how to solve this problem

OpenStudy (fibonaccichick666):

so, how do you calculate slope?

OpenStudy (j.f):

slope=rise/run

OpenStudy (nincompoop):

what does that mean?

OpenStudy (fibonaccichick666):

like can you explicitly state the formula

OpenStudy (fibonaccichick666):

@J.F

OpenStudy (j.f):

y=mx+b, m being slope and b being y intercept value

OpenStudy (nincompoop):

\(\large slope = m = \frac{y_2 - y_1}{x_2-x_1} \rightarrow 3 = \frac{a-5}{4-3} \)

OpenStudy (nincompoop):

solve for a

OpenStudy (fibonaccichick666):

so, if you have two points, how do you calculate slope?

OpenStudy (nincompoop):



OpenStudy (j.f):

nin, a=8?

OpenStudy (nincompoop):

solve it

OpenStudy (nincompoop):

show how you got 8

OpenStudy (fibonaccichick666):

I agree with nin, please show your work

OpenStudy (j.f):

3/1=a-5/4-3 3(4-3)=1(a-5)= 12-9=a-5 5+12-9=a 8=a

OpenStudy (fibonaccichick666):

you could have done 4-3 first, but that's ok

OpenStudy (j.f):

So did I do something wrong?

OpenStudy (fibonaccichick666):

no, but you could have made your life a little eaiser

OpenStudy (fibonaccichick666):

, it is correct nonetheless

OpenStudy (j.f):

yeah, I was taught horribly in math. I kinda learned things weird. But its what I know

OpenStudy (fibonaccichick666):

hey, it's all legal, and it works. I'd just say between steps ask yourself, can I simplify anything? combine any like terms? Then go

OpenStudy (nincompoop):

what she is saying is that if you can simplify things early, do it

OpenStudy (fibonaccichick666):

BUt you should really memorize the slope formula. I tis REALLY REALLY useful

OpenStudy (nincompoop):

I honestly didn't memorize anything except areas until I started college, but I was taught concepts outside of the formulas so it was all good.

OpenStudy (fibonaccichick666):

well, we could derive slope, but if you don't like derivations, you're better off memorizing it

OpenStudy (nincompoop):

if you think of slope, you think about the steepness - how sharp/fast it goes up or down for every time you go forward

OpenStudy (j.f):

yeah, They say the slope method is crucial to understanding Algebra. I on the other hand just suck at remembering math principles. Its not that I don't know math, I just forget suuper easy. Probably from lack of repetition

OpenStudy (fibonaccichick666):

oh, then shall nin and I teach you how to derive it?

OpenStudy (nincompoop):